A hockey puck (mass = 2.5 kg) leaves the players stick with a speed of 25 m/s and slides on the ice for 90 meters before coming to rest.

A) What is the magnitude of the acceleration on the puck?
m/s2

Tries 0/2


B) What is the magnitude of the friction force exerted on the puck due to the ice?
N

Tries 0/2


C) What is the normal force on the puck?
N

Tries 0/2


D) What is the friction coefficient between the puck and the ice?
(unitless)
Answer #1

a)

vo = initial speed = 25 m/s

v = final speed = 0 m/s (since puck comes to a stop )

x = distance traveled before coming to stop = 90 m

a = acceleration of puck

Using the kinematics equation

v2 = vo2 + 2 a x

02 = 252 + 2 a (90)

0 = 625 + 180 a

a = - 625/180

a = - 3.5 m/s2

magnitude of acceleration : 3.5 m/s2

b)

f = magnitude of static friction

magnitude of static friction is given as

f = ma

f = (2.5) (3.5)

f = 8.75 N

c)

Normal force on the puck is given as

N = mg

N = (2.5) (9.8)

N = 24.5 N

d)

Coefficient of friction is given as

\mu = f/N

\mu = 8.75/24.5

\mu = 0.36